OpenStudy (anonymous):

Widget Wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $23 to produce 2 widgets, $103 to produce 4 widgets, and $631 to produce 10 widgets. find the cost of producing 6 widgets.

OpenStudy (anonymous):

NEED HELP !! D;

OpenStudy (phi):

this looks painful. they are saying c(x) = a x^2 + b x + d and are given (2,23) , (4,103) and (10,631) are points on this parabola.

OpenStudy (anonymous):

yea it is painful

OpenStudy (anonymous):

do you want to see a picture if thatll help?

OpenStudy (phi):

to find c(6) we need to find the equation , which means find a, b and d to do that we use the data they gave us. using (4,103) we write 103= 16a + 4b + d using (2,23) we write 23 = 4a + 2b + d subtract these 2 equations 103= 16a + 4b + d 23 = 4a + 2b + d ------------------ 80 = 12a + 2b simplify (divide all terms by 2) 40 = 6a + b now do the same thing with (10,631) and (2,23)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im confused on how you got 16

OpenStudy (phi):

using (4,103) in c(x) = a x^2 + b x + d replace c(x) with 103 and x with 4: 103 = a (4^2) + b*4 + d 4*4 is 16. we get 103 = 16a + 4 b + d (we usually write the letters after the numbers...easier to read)

OpenStudy (phi):

if they give a graph, perhaps we can read off the answer ?

OpenStudy (anonymous):

theres no graph =l

OpenStudy (anonymous):

631=20a+10b +d . is that right?

OpenStudy (phi):

20a? remember is is x^2 which means x*x (not 2*x, which is different)

OpenStudy (anonymous):

oh 100a+10b+d

OpenStudy (phi):

yes 631= 100a+10b+d now use another equation, example: 23 = 4a + 2b + d (from (2,23) point) 631= 100a+10b+d 23 = 4a + 2b + d subtract the 2nd equation from the first

OpenStudy (anonymous):

608=96a+8b+d ?

OpenStudy (phi):

d-d is 0

OpenStudy (anonymous):

okay just 608=96a+8b

OpenStudy (phi):

and make our problem easier... divide all terms by 8

OpenStudy (anonymous):

76=96+8

OpenStudy (anonymous):

i dont know what to do after that

OpenStudy (phi):

the letters do not disappear. Example a*8/8 = a * 8/8 = a*1= a 96a/8 = 96/8 * a = 12 a try again

OpenStudy (anonymous):

wait where did you get 12

OpenStudy (anonymous):

76=96a+8b

OpenStudy (phi):

you start with 608=96a+8b you can divide both sides by 8 (and things stay equal) \[ \frac{608}{8} = \frac{96a+8b}{8} \] on the right side we can write it as two fractions because \[ \frac{96a}{8}+\frac{8b}{8} = \frac{96a+8b}{8} \] \[ \frac{608}{8} = \frac{96a}{8}+\frac{8b}{8} \] now simplify (and the letters do not disappear)

OpenStudy (phi):

to make it more clear, it is the same as \[ \frac{608}{8} = \frac{96}{8}a+\frac{8}{8}b \]

OpenStudy (anonymous):

76=12a+1b

OpenStudy (phi):

yes. we now have two equations with two unknowns. From up above we have 40 = 6a + b and your equation 76=12a+ b (we can write 1b as just b) any idea what to do?

OpenStudy (anonymous):

uhh. minus it? 36=6a+1b

OpenStudy (anonymous):

i have no clue sorry

OpenStudy (phi):

yes that is the right idea. But do it correctly (carefully)

OpenStudy (anonymous):

36=6a+b

OpenStudy (phi):

I guess you did 76-40 =12a-6a + b - b ?

OpenStudy (anonymous):

umm no

OpenStudy (anonymous):

omg nvm i read that wrong

OpenStudy (phi):

when you simplify the b's go away

OpenStudy (anonymous):

oh yeah i did

OpenStudy (anonymous):

just 36=6a then?

OpenStudy (phi):

yes. In fact when you see a single b in both equations you should think "aha! subtraction will make the b's go away"

OpenStudy (phi):

now you have 6a= 36 solve for a (divide both sides by 6)

OpenStudy (anonymous):

uhh i just have 6

OpenStudy (phi):

6a/6 = 36/6 a= 6

OpenStudy (anonymous):

but the problem asks the total cost for 6 widgets

OpenStudy (phi):

yes, but this is a long problem. however, we are close. remember what we are doing. they told us we have a quadratic which means we have something that looks like \[c(x) = a x^2 + b x + d \] we found a=6 we need to find b and d but we know 40 = 6a + b and 76=12a+1b use either of those equations (with a replaced by 6) to find b can you do that ?

OpenStudy (anonymous):

i dont think so

OpenStudy (phi):

which equation do you want to use ?

OpenStudy (anonymous):

17=12a+1b

OpenStudy (anonymous):

i mean 76

OpenStudy (phi):

ok (because using 17 would not work) 76=12a+1b we found a=6 erase the a and put in 6 in its place. and remember the equation means 76= 12*a + 1*b

OpenStudy (anonymous):

76=72a?

OpenStudy (phi):

read this carefully: *erase the a* (no "a"!) and put in 6 in its place and leave everything else as is. 76= 12*a + 1*b

OpenStudy (anonymous):

76=72 ..?

OpenStudy (phi):

you are "doing too much" erase the "a" 76 = 12 * ___ + 1*b put 6 in its place 76= 12 * 6 + 1*b

OpenStudy (anonymous):

okay

OpenStudy (phi):

now we can do 12*6 = 72 76 = 72 + 1b add -72 to both sides

OpenStudy (anonymous):

148=144

OpenStudy (phi):

do it in smaller steps. just write -72 on both sides of the equation 76 = 72 + 1b

OpenStudy (anonymous):

sorry i dont have much time

OpenStudy (anonymous):

-72 76= 72 + 1b -72

OpenStudy (anonymous):

?

OpenStudy (phi):

yes but we are adding so it should look like this -72 + 76= 72 + 1b -72 simplify both sides

OpenStudy (phi):

76-72 on the left side 72-72 on the right side you get 4 = 0+1b or just 4 = b

OpenStudy (anonymous):

okay

OpenStudy (phi):

now find d. use 23 = 4a + 2b + d remember a=6 and b=4

OpenStudy (anonymous):

23=24a+8b ?

OpenStudy (phi):

no, 23 = 4*6 + 2*4 + d

OpenStudy (phi):

23= 24+8 + d 23= 32 + d add -32 to both sides 23 - 32 = 32 - 32 + d

OpenStudy (anonymous):

55=64+d

OpenStudy (phi):

your calculator is broken. 23-32 is -9 and 32 - 32 is 0

OpenStudy (phi):

you should get d= -9 now you have c(x) = 6x^2 + 4x - 9 find c(6) which means replace x with 6 and simplify

OpenStudy (anonymous):

okay umm

OpenStudy (phi):

in other words figure out 6* 6^2 + 4*6 - 9

OpenStudy (anonymous):

i got 231

OpenStudy (anonymous):

231+29-9

OpenStudy (anonymous):

is 231 correct?

OpenStudy (anonymous):

omg it was! thank u soo much

OpenStudy (anonymous):

sorry for probably giving you a headache

OpenStudy (phi):

yes. it was a tough problem